OpenStudy (anonymous):

Find f(-7) if f(x) = 2x + 8. How do i do this?

OpenStudy (cggurumanjunath):

put x =-7 in 2x+8

OpenStudy (anonymous):

Substitute for x with -7 so it will be : f(x) = 2x + 8 f(-7)=2(-7)+8 f(-7)=-14+8=-6

OpenStudy (cggurumanjunath):

2(-7)+8

OpenStudy (anonymous):

thanks guys i got it now

OpenStudy (cggurumanjunath):

:)